OpenStudy (anonymous):

112 in feet and inches

OpenStudy (anonymous):

112 what?

OpenStudy (anonymous):

in= inches

OpenStudy (anonymous):

112iches converted to feet and inches

OpenStudy (anonymous):

112 / 12 = 9.333 which equal 9 feet 4 inches

OpenStudy (anonymous):

thx

OpenStudy (doc.brown):

Better yet.\[112in\times\frac{ft}{12in}=9ft...\]\[9ft\times\frac{12in}{ft}=108in\]\[112in-108in=4in\]

OpenStudy (anonymous):

Thanks for explaining :P

OpenStudy (doc.brown):

:P